OpenStudy (anonymous):

How many possibel values of k if equation kx^2-(k+2)x+k=0 and x^2-(k+1)x+k=0 has one same solution? Find the sum of the integer values of k. PLEASE HELP IM IN 6TH GRADE IM INNOCENT AND SHOULDN'T BE DOING THESE PROBLEMS!!

OpenStudy (anonymous):

Anyone? IF you answer and get it right i will fan you and give ou a medal

OpenStudy (anonymous):

\[K=2x/x ^{2}-x+1\]

OpenStudy (anonymous):

i need a value. answer the second quetion plz

OpenStudy (anonymous):

i fanned both of u

OpenStudy (anonymous):

how did u get what k equals? thx i think it is right

OpenStudy (anonymous):

It's not. Try using the quadratic formula.

OpenStudy (anonymous):

i did!

OpenStudy (anonymous):

it didn't work, i didnt get the sum of the integer values of k

OpenStudy (anonymous):

they have to have one ame answet

OpenStudy (anonymous):

fanned Sithsandgiggles

OpenStudy (anonymous):

fanned jonask

OpenStudy (anonymous):

fanned anyone who saw my question

OpenStudy (anonymous):

thx everyone i appreciate it

OpenStudy (anonymous):

\(kx^2-(k+2)x+k=0\) use the quadratic formular \[\dfrac{-b\pm sqrt{b^2-4ac}}{2a}\] for first one \[x=\frac{k+2\pm \sqrt{(k+2)^2-4k^2}}{2k}\] \(x^2-(k+1)x+k=0\) second one \[x=\frac{k+1\pm \sqrt{(k+1)^2-4k}}{2}\] you have to make this two solutions equal

OpenStudy (anonymous):

gues that makes sense. but how do u make them equal? there are two variables and you can't must plug in variables. there could be millions of answers that would work.

OpenStudy (anonymous):

i already tried that

OpenStudy (anonymous):

notice that there are two solutions from each of those,and you have four equations,you will decide wich pairs are equal

OpenStudy (anonymous):

if they are not equal ,you will get no solution for the equality

OpenStudy (anonymous):

four equations, so how do u know which 2 are equal?

OpenStudy (anonymous):

there is a solution, i just dont know what it is. I have to find the sum of all the values of k

OpenStudy (anonymous):

the easy way is to think like this:for equations to have same roots,the need to satisfy the following conditions Sum of roots=\(\dfrac{-b}{a}\) product of roots =\(\dfrac{c}{a}\)

OpenStudy (anonymous):

ya, thats the vieta theroem. but how does that help me?

OpenStudy (anonymous):

you could use the LHS from the first solution and RHS from the second equation

OpenStudy (anonymous):

whats a lhs and a rhs

OpenStudy (anonymous):

from first equation \[x_1+x_2=\frac{k+2+ \sqrt{(k+2)^2-4k^2}}{2k}+\frac{k+2- \sqrt{(k+2)^2-4k^2}}{2k}=\dfrac{k+2}{k}\] \[=-\frac{b}{a}_{from \space 2nd \space equation}=\frac{k+1}{2}\]

OpenStudy (anonymous):

LHS=left hand side RHS =Right hand side

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

omg i get it now thx

OpenStudy (anonymous):

i'm glad you do,you show understanding i wish you'll be able to solve it

OpenStudy (anonymous):

\[\frac{k+2}{k}=\frac{k+1}{2}\]